# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ E.2 Proofs of Lemmas for Lemma 1
$\leq(c_{\mu}+1)\left(\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}\right)+(c_{\mu}+1)\left\{\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right\}^{2}$
@@
$=(c_{\mu}+1)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left\{\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right\}^{2}\right),$ (E.9)
where the second inequality holds due to the AM-GM inequality. Finally, combining (E.6) and (E.9), by setting $\mu=1/t$, we have
@@
E.2.3 Proof of Lemma E.3
Proof of Lemma E.3.
@@
$\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq-\log\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[e^{-\ell_{s}(\mathbf{w})}\right]\right)=-\log\left(\frac{1}{Z_{s}}\int_{\mathbb{R}^{d}}e^{-L_{s}(\mathbf{w})}\mathrm{d}\mathbf{w}\right),$ (E.10)
@@
$\tilde{L}_{s}(\mathbf{w})=L_{s}(\mathbf{w}_{s+1})+\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}.$ (E.11)
Applying Lemma E.8 and considering the fact that $\ell_{s}$ is $3\sqrt{2}$-self-concordant-like function by Proposition B.1, we have
$L_{s}(\mathbf{w})\leq\tilde{L}_{s}(\mathbf{w})+e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla\ell_{s}(\mathbf{w}_{s+1})}^{2}.$ (E.12)
@@
Then, we can establish a lower bound for the expectation in (E.10) as follows:
@@
$=\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))}{Z_{s}}\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\cdot\exp(-\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle)\mathrm{d}\mathbf{w},$ (E.13)
where the first inequality holds by (E.12) and the last equality holds by (E.11). We define $\tilde{Z}_{s+1}=\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\mathrm{d}\mathbf{w}\leq+\infty$. Moreover, we denote the distribution whose density function is $\tilde{f}_{s+1}(\mathbf{w})/\tilde{Z}_{s+1}$ as $\tilde{P}_{s+1}$. Then, we can rewrite the equation (E.13) as follows:
@@
$=\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))\tilde{Z}_{s+1}}{Z_{s}},$ (E.14)
where the second inequality follows from Jensen’s inequality, and the last inequality holds because $\tilde{P}_{s+1}$ is symmetric around $\mathbf{w}_{s+1}$, thus $\mathbb{E}_{\mathbf{w}\sim\tilde{P}_{s+1}}\left[\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle\right]=0$.
By plugging (E.14) into (E.10), we have
$\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq L_{s}(\mathbf{w}_{s+1})+\log Z_{s}-\log\tilde{Z}_{s+1}.$ (E.15)
@@
$=-\log Z_{s}+\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right],$ (E.16)
where $\widehat{P}_{s+1}=\mathcal{N}(\mathbf{w}_{s+1},cH_{s}^{-1})$ and $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}$. In (E.16), the inequality holds due to Jensen’s inequality, and the last inequality is by the fact that $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}=\sqrt{(2\pi)^{d}c|H_{s}^{-1}|}=Z_{s}$.
By applying the Cauchy-Schwarz inequality, we can further bound the second term on the right-hand side of (E.16) by
@@
$\leq\underbrace{\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}}_{\texttt{(a)-1}}\underbrace{\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}}_{\texttt{(a)-2}}.$ (E.17)
@@
$\sum_{j=1}^{d}\sqrt{c\lambda_{j}\left(H_{s}^{-1}\right)}X_{j}\mathbf{e}_{j},\quad\text{where }X_{j}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,1),\forall j\in[d],$ (E.18)
and $\lambda_{j}\left(H_{s}^{-1}\right)$ denotes the $j$-th largest eigenvalue of $H_{s}^{-1}$. Then, we can bound the term (a)-1 in (E.17) as follows:
@@
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}\leq\mathbb{E}_{X\sim\chi^{2}}\left[e^{\frac{X}{4}}\right]\leq\sqrt{2},$ (E.19)
@@
where $\bar{H}_{s}=(\nabla^{2}\ell_{s}(\mathbf{w}_{s+1}))^{-1/2}H_{s}(\nabla^{2}\ell_{s}(\mathbf{w}_{s+1}))^{-1/2}$. Let $\bar{\lambda}_{j}=\lambda_{j}\left(c\bar{H}_{s}^{-1}\right)$ be the $j$-th largest eigenvalue of the matrix $c\bar{H}_{s}^{-1}$. Then, conducting an analysis similar to that in equation (E.18) yields that
@@
where the last inequality holds due to $\mathbb{E}_{X_{j},X_{j^{\prime}}\sim\mathcal{N}(0,1)}[X_{j}^{2}X_{j^{\prime}}^{2}]\leq 3$ for all $j,j^{\prime}\in[d]$, and the last equality holds because $\sum_{j}^{d}\bar{\lambda}_{j}=\operatorname{Tr}\left(c\bar{H}_{s}^{-1}\right)$. Here, $\operatorname{Tr}(A)$ denotes the trace of the matrix $A$.
@@
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}\leq\sqrt{3}c\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}.$ (E.20)
By plugging (E.19) and (E.20) into (E.17), we have
$\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right]\leq\sqrt{6}c\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}.$ (E.21)
Combining (E.15), (E.16), and (E.21), and taking summation over $s$, we derive that
@@
E.3 Technical Lemmas for Lemma 1
Lemma E.4 (Proposition 4.1 of Campolongo and Orabona 7).
@@ E.3 Technical Lemmas for Lemma 1
Lemma E.5 (Lemma 15 of Zhang and Sugiyama 29).
@@
Lemma E.6 (Lemma 1 of Zhang and Sugiyama 29).
@@
Lemma E.7 (Lemma 17 of Zhang and Sugiyama 29).
@@
Lemma E.8 (Lemma 18 of Zhang and Sugiyama 29).
@@
Appendix F Proofs of Theorem 3
In this section, we provide the proof of Theorem 3. In addition to the adversarial construction presented in Section C.1, we construct the non-uniform rewards.
F.1 Adversarial Rewards Construction
Under the adversarial construction in Section C.1, we observe that there are $K$ identical context vectors, invariant across rounds $t$. Therefore, in total, there are $N=K\cdot\binom{d}{d/4}$ items. Let the rewards be also time-invariant. Given $\mathbf{w}_{V}$, we define a unique item $i^{\star}\in[N]$ as an item that maximizes $x_{i}^{\top}\mathbf{w}_{V}$, i.e., $x_{i^{\star}}=x_{V}$, and has a reward of $1$, i.e., $r_{i^{\star}}=1$. Then, we construct the non-uniform rewards as follows:
$r_{i}=\begin{cases}1,&\text{for }i=i^{\star}\\ \gamma,&\text{for }i\neq i^{\star},\end{cases}$ (F.1)
@@ F.1 Adversarial Rewards Construction
F.2 Main Proof of Theorem 3
@@
Given the rewards construction as (F.1), we can bound the expected revenue for any assortment as follows:
Lemma F.1.
Under the same parameters and context vectors as those in Section C, if the rewards are constructed according to Equation (F.1), for any $S\in\mathcal{S}$, we have
Furthermore, the following lemma shows that the rewards in the optimal assortment $S_{t}^{\star}$ are larger than their expected revenues.
Lemma F.2.
@@
Lemma F.2 implies that $S_{t}^{\star}$ contains only one item $i^{\star}$ because $\gamma\leq R(S^{\star},\mathbf{w}_{v})$.
@@
where the first equality holds by Lemma F.1, and the inequality holds by Lemma F.2. Hence, the problem is not easier than solving the MNL bandit problems with the assortment size $1$, i.e., $K=1$. By putting $K=1$ and $v_{0}=\Theta(1)$ in Theorem 1, we derive that
@@
F.3 Proofs of Lemmas for Theorem 3
F.3.1 Proof of Lemma F.1
Proof of Lemma F.1.

### v6

@@ F.2 Proofs of Lemmas for Lemma 1
$\leq(c_{\mu}+1)\left(\lambda+\sum_{s=1}^{t}\|\mathbf{d}_{s}\|_{\nabla\bm{\sigma}_{s}(\mathbf{z}_{s}^{\star})}^{2}\right)+(c_{\mu}+1)\left(\frac{\sqrt{\lambda}}{4}+4\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}$
@@
$=(c_{\mu}+1)\left(\frac{17}{16}\lambda+2\sqrt{\lambda}\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)+16\left(\log\left(\frac{2\sqrt{1+2t}}{\delta}\right)\right)^{2}\right),$ (F.9)
where the third inequality holds due to the AM-GM inequality. Finally, combining (F.6) and (F.9), by setting $\mu=1/t$, we have
@@
F.2.3 Proof of Lemma F.3
Proof of Lemma F.3.
@@
$\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq-\log\left(\mathbb{E}_{\mathbf{w}\sim P_{s}}\left[e^{-\ell_{s}(\mathbf{w})}\right]\right)=-\log\left(\frac{1}{Z_{s}}\int_{\mathbb{R}^{d}}e^{-L_{s}(\mathbf{w})}\mathrm{d}\mathbf{w}\right),$ (F.10)
@@
$\tilde{L}_{s}(\mathbf{w})=L_{s}(\mathbf{w}_{s+1})+\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}.$ (F.11)
Applying Lemma F.8 and considering the fact that $\ell_{s}$ is $3\sqrt{2}$-self-concordant-like function by Proposition C.1, we have
$L_{s}(\mathbf{w})\leq\tilde{L}_{s}(\mathbf{w})+e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla\ell_{s}(\mathbf{w}_{s+1})}^{2}.$ (F.12)
@@
Then, we can establish a lower bound for the expectation in (F.10) as follows:
@@
$=\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))}{Z_{s}}\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\cdot\exp(-\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle)\mathrm{d}\mathbf{w},$ (F.13)
where the first inequality holds by (F.12) and the last equality holds by (F.11). We define $\tilde{Z}_{s+1}=\int_{\mathbb{R}^{d}}\tilde{f}_{s+1}(\mathbf{w})\mathrm{d}\mathbf{w}\leq+\infty$. Moreover, we denote the distribution whose density function is $\tilde{f}_{s+1}(\mathbf{w})/\tilde{Z}_{s+1}$ as $\tilde{P}_{s+1}$. Then, we can rewrite the equation (F.13) as follows:
@@
$=\frac{\exp(-L_{s}(\mathbf{w}_{s+1}))\tilde{Z}_{s+1}}{Z_{s}},$ (F.14)
where the second inequality follows from Jensen’s inequality, and the equality holds because $\tilde{P}_{s+1}$ is symmetric around $\mathbf{w}_{s+1}$, thus $\mathbb{E}_{\mathbf{w}\sim\tilde{P}_{s+1}}\left[\langle\nabla L_{s}(\mathbf{w}_{s+1}),\mathbf{w}-\mathbf{w}_{s+1}\rangle\right]=0$.
By plugging (F.14) into (F.10), we have
$\ell(\tilde{\mathbf{z}}_{s},\mathbf{y}_{s})\leq L_{s}(\mathbf{w}_{s+1})+\log Z_{s}-\log\tilde{Z}_{s+1}.$ (F.15)
@@
$=-\log Z_{s}+\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right],$ (F.16)
where $\widehat{P}_{s+1}=\mathcal{N}(\mathbf{w}_{s+1},cH_{s}^{-1})$ and $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}$. In (F.16), the inequality holds due to Jensen’s inequality, and the last inequality is by the fact that $\widehat{Z}_{s+1}=\int_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{H_{s}}^{2}\right)\mathrm{d}\mathbf{w}=\sqrt{(2\pi)^{d}c|H_{s}^{-1}|}=Z_{s}$.
By applying the Cauchy-Schwarz inequality, we can further bound the second term on the right-hand side of (F.16) by
@@
$\leq\underbrace{\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}}_{\texttt{(a)-1}}\underbrace{\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}}_{\texttt{(a)-2}}.$ (F.17)
@@
$\sum_{j=1}^{d}\sqrt{c\lambda_{j}\left(H_{s}^{-1}\right)}X_{j}\mathbf{e}_{j},\quad\text{where }X_{j}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,1),\forall j\in[d],$ (F.18)
and $\lambda_{j}\left(H_{s}^{-1}\right)$ denotes the $j$-th largest eigenvalue of $H_{s}^{-1}$. Then, we can bound the term (a)-1 in (F.17) as follows:
@@
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{36\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\right]}\leq\mathbb{E}_{X\sim\chi^{2}}\left[e^{\frac{X}{4}}\right]\leq\sqrt{2},$ (F.19)
@@
where $\bar{H}_{s}=(\nabla^{2}\ell_{s}(\mathbf{w}_{s+1}))^{-1/2}H_{s}(\nabla^{2}\ell_{s}(\mathbf{w}_{s+1}))^{-1/2}$. Let $\bar{\lambda}_{j}=\lambda_{j}\left(c\bar{H}_{s}^{-1}\right)$ be the $j$-th largest eigenvalue of the matrix $c\bar{H}_{s}^{-1}$. Then, conducting an analysis similar to that in equation (F.18) yields that
@@
where the inequality holds due to $\mathbb{E}_{X_{j},X_{j^{\prime}}\sim\mathcal{N}(0,1)}[X_{j}^{2}X_{j^{\prime}}^{2}]\leq 3$ for all $j,j^{\prime}\in[d]$, and the last equality holds because $\sum_{j}^{d}\bar{\lambda}_{j}=\operatorname{Tr}\left(c\bar{H}_{s}^{-1}\right)$. Here, $\operatorname{Tr}(A)$ denotes the trace of the matrix $A$.
@@
$\sqrt{\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{4}\right]}\leq\sqrt{3}c\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}.$ (F.20)
By plugging (F.19) and (F.20) into (F.17), we have
$\mathbb{E}_{\mathbf{w}\sim\widehat{P}_{s+1}}\left[e^{18\|\mathbf{w}-\mathbf{w}_{s+1}\|_{2}^{2}}\|\mathbf{w}-\mathbf{w}_{s+1}\|_{\nabla^{2}\ell_{s}(\mathbf{w}_{s+1})}^{2}\right]\leq\sqrt{6}c\log\frac{\operatorname{det}(M_{s+1})}{\operatorname{det}(M_{s})}.$ (F.21)
Combining (F.15), (F.16), and (F.21), and taking summation over $s$, we derive that
@@
F.3 Technical Lemmas for Lemma 1
Lemma F.4 (Proposition 4.1 of Campolongo and Orabona 10).
@@ F.3 Technical Lemmas for Lemma 1
Lemma F.5 (Lemma 15 of Zhang and Sugiyama 52).
@@
Lemma F.6 (Lemma 1 of Zhang and Sugiyama 52).
@@
Lemma F.7 (Lemma 17 of Zhang and Sugiyama 52).
@@
Lemma F.8 (Lemma 18 of Zhang and Sugiyama 52).
@@
Appendix G Proofs of Theorem 3
In this section, we provide the proof of Theorem 3. In addition to the adversarial construction presented in Section D.1, we construct the adversarial non-uniform rewards.
G.1 Adversarial Rewards Construction
Under the adversarial construction in Section D.1, we observe that there are $K$ identical context vectors, invariant across rounds $t$. Therefore, in total, there are $N=K\cdot\binom{d}{d/4}$ items. Let the rewards be also time-invariant. Given $\mathbf{w}_{V}$, we define a unique item $i^{\star}\in[N]$ as an item that maximizes $x_{i}^{\top}\mathbf{w}_{V}$, i.e., $x_{i^{\star}}=x_{V}$, and has a reward of $1$, i.e., $r_{i^{\star}}=1$. Then, we construct the non-uniform rewards as follows:
$r_{i}=\begin{cases}1,&\text{for }i=i^{\star}\\ \gamma,&\text{for }i\neq i^{\star},\end{cases}$ (G.1)
@@ G.1 Adversarial Rewards Construction
G.2 Main Proof of Theorem 3
@@
Given the rewards construction as (G.1), any reward in the optimal assortment $S_{t}^{\star}$ is larger than the expected revenues.
Lemma G.1.
@@
Lemma G.1 implies that $S^{\star}$ contains only one item $i^{\star}$. This is because if $S^{\star}=\{x_{i^{\star}}\}$, adding any item $i\neq i^{\star}$ to the assortment results in lower expected revenue, since $r_{i}=\gamma\leq R(S^{\star}=\{x_{i^{\star}}\},\mathbf{w}_{v})$.
Furthermore, we can bound the expected revenue for any assortment as follows:
Lemma G.2.
Under the same parameters and context vectors as those in Section D, if the rewards are constructed according to Equation (G.1), for any $S\in\mathcal{S}$, we have
@@
where the first equality holds becauset $S^{\star}$ contains only one item $i^{\star}$ by Lemma G.1 (and recall that $x_{i^{\star}}=x_{V}$), and the inequality holds by Lemma G.2. Hence, the problem is not easier than solving the MNL bandit problems with the assortment size $1$, i.e., $K=1$. By putting $K=1$ and $v_{0}=\Theta(1)$ in Theorem 1, we derive that
@@
G.3 Proofs of Lemmas for Theorem 3
G.3.1 Proof of Lemma G.1
Proof of Lemma G.1.
We prove by contradiction. Assume that there exists $i\in S^{\star}$ such that $r_{i}<R(S^{\star},\mathbf{w}_{V})$. Then, removing the item $i$ from the assortment $S^{\star}$ yields higher expected revenue. This contradicts the optimality of $S^{\star}$. Thus, we have
$r_{i}\geq R(S^{\star},\mathbf{w}_{V}),\quad\forall i\in S^{\star}.$
This concludes the proof. ∎
G.3.2 Proof of Lemma G.2
Proof of Lemma G.2.